OpenStudy (anonymous):

Complete the blank 17+2=(12- ___ )+8___

OpenStudy (anonymous):

ok. so 19 = (12 - x ) + 8x what number can we use to fit this equation? first try 2: (12 - 2) + 8(2) = 10 + 16 = 26 too many. let's try 1: (12 - 1) + 8(1) = 11 + 8 = 19 :) the blank should be 1